The reaction X2(g)<->2X(g) occurs in a closed reaction vessel at constant volume and temperature. Initially, the vessel contains only X2 at a pressure of 1.75 atm. After the reaction reaches equilibrium, the total pressure is 2.65atm.
what is the value of the equilibrium constant, Kp, for the reaction?
Answer #1

X2(g)   <->   2X(g)
1.75           0   (initial)
1.75-x       2x   (equilibrium)

Total pressure at equilibrium = 1.75-x+2x
2.65 = 1.75+x
x = 0.90 atm

Now use:
Kp = p(X)^2 / p(X2)
= (2x)^2 / (1.75-x)
= (2*0.90)^2 / (1.75 - 0.90)
= 3.24 / 0.85
= 3.81

Answer: 3.81